OpenStudy (anonymous):

if sinx= 12/13 find sec

OpenStudy (anonymous):

that csc

OpenStudy (lgbasallote):

lol sorry...

OpenStudy (chaise):

Are you aware that: \[\sin^2\theta+\cos^2\theta=1\]

OpenStudy (anonymous):

yes

OpenStudy (chaise):

Perhaps you could utilise this identity in solving your answer :)

OpenStudy (lgbasallote):

well find your third side by doing \[\sqrt{13^2 - 12^2}\]

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow sinX = {opposite \over hypotenuse} }\) Clearly, opposite is 12 and 13 is hypotenuse. \(\Large \color{MidnightBlue}{\Rightarrow secX = {hypotenuse \over adjacent} }\) Hypotenuse is 13, of course. Now, we know that 13 is involved in Pythagorean triplet 5,12,13 where 13 is the hypotenuse. Clearly, 12 is opposite so the only side left is 5(adjacent) \(\Large \color{MidnightBlue}{\Rightarrow secX = {13 \over 5} = 2{3 \over 5} }\)

Parth (parthkohli):

Was I right?

OpenStudy (anonymous):

@chaise how would i use it

Parth (parthkohli):

@viictor As far as I am concerned, that has no use. See my post please.

OpenStudy (chaise):

Parth is probably right ^

Parth (parthkohli):

BTW, you needed csc or sec?

Parth (parthkohli):

If you needed csc, then: \(\Large \color{MidnightBlue}{\Rightarrow cscX = {1 \over sinX} }\)

OpenStudy (anonymous):

@chaise i believe youre right

Parth (parthkohli):

Or, you can say that csc is the reciprocal of sin.

OpenStudy (anonymous):

i need sec

Parth (parthkohli):

Then I have given the value of secX.

OpenStudy (lgbasallote):

for all that matters....kohli is right

OpenStudy (anonymous):

im not using a triangle!

Parth (parthkohli):

Trig is all about triangles! :P

OpenStudy (lgbasallote):

where there is sine..there is automatically a triangle young padawan

OpenStudy (kropot72):

@viictor BTW why did your second posting say you wanted cosec?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow Trigonometry = \text{Measure of triangles.} }\)

Parth (parthkohli):

@kropot72 I've already mentioned csc.

OpenStudy (lgbasallote):

well goodluck finding someone who can solve it by sin^2 + cos^2 = 1

OpenStudy (kropot72):

That's what I say too :0